SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Oct. 11, 2005)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Fang Fang, Selina Li, Dan Walker.

Operations

  • PEP was able to reach a peak lumi of just over 1x10^34 last Sunday. The October shutdown started on Monday Oct.10. We started to test the new TSF firmware right after the beam dump on Monday. See the DCT/DCZ section below for more detailed info.

EMT

  • The new EMT online code will be published and be in the online release.
  • Dan will also tag L1TFMon that includes the EMT turn-on curves.
  • Last Friday we got paged about the EMT fastmon plots in JAS for being different from the reference plot in one of runs that have high background. Rainer suggested to find out if it was trickle injection related.

DCT/DCZ

  • We tested Marc's new TSF firmware at both the test stand and IR2. See the summary here. The conclusion is that it works very well, except that we still need to understand the bit corruption problem that found out from the ROM logfiles for both the TSF-X and TSF-Y. Currently, we don't know if it's caused by the firmware or by the fact that some of the boards might not be inserted properly. This is under investigation. Rainer suggested to run L1TPlayback to find any discrepancy like this. Fang will look at the code for L1DctNewTsfL1Accept which prints out these error messages such as "unphysical sl" and "bit corrupt", and modify it to provide us more information about this problem. Rainer asked where the TSF damages come from. We need to understand why there are inconsistent data from the boards.

L3

  • As the lumi gets higher, it might be necessary to watch the rate and re-think about prescaling bhabha etc.
  • PEP will be pushing for aggressive injection when we start taking data again in November. If it happens, we want a two-dimentional time filter to help monitoring. Last Sunday when we had highest luminosity, the rate was close to the goal. At that time, Al thinks that PEP only pushed for HER not LER.
  • Al suggested that we need a revised trickle injetion task force. We need to revise the trickle injection monitoring plots, and dispaly the distribution of the Weaver plot. The question is who would do it? PEP also expressed interest in seeing the trickle injection plots.
  • Jim Hamilton finally have core dump for L3 created when there is a node trim so that we can study what caused the trim. This revealed a bug in the code for which L3 was waiting for trickle to finish in some of the node trims happened in the past. This is more like the OEP problem than L3. It explains some of the node trims but not all.

Data Quality and Monitoring

  • The threshold for notifying a TRG expert about the EMT/DCT->GLT transmission error has been raised to 60 rather than 30 to take into account the fact that we have higher lumi now.
  • Kris is trying to get Elf working now to process run 58747 for the TSF firmware test.
  • L3 tracking efficiency seems to be lower for SP8 in run 5. But Rainer seemed to remember that this has always been an issue. Kris will be signing off the SP8 validation soon.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.